OpenStudy (anonymous):

A 20-cm long and 1.0-mm diameter thick copper wire (resistivity 1.7 × 10-8 Ωm) is connected across a 3-V battery. The current through the wire is 1-7 A. 2-125 A. 3-693 mA. 4-693 A. 5-789 A.

OpenStudy (anonymous):

@experimentx @jhonyy9 @ishaan94 who know the answer gyus but please quickly i have 15 min only to solve these one :( :\

OpenStudy (anonymous):

@siddharth18 do u can solve it :)

OpenStudy (anonymous):

from ohm's we know that v=ir now we have v=3V r=(resistivity*length)/area of cross section of the wire i(current)=v/r radius of the wire=0.5mm now u have all the values,just substitute them in the formula for i and calculate

OpenStudy (anonymous):

area of cross section is given by pi*r^2

OpenStudy (anonymous):

make sure u use SI units

OpenStudy (anonymous):

are u sure from how to get the area ?

OpenStudy (anonymous):

yes i m

OpenStudy (anonymous):

but i get the answer 0.01298 can u check and tell me what u get please ?

OpenStudy (anonymous):

but i get the answer 0.01298 can u @siddharth18 check and tell me what u get please ?

OpenStudy (anonymous):

my answer is 789amperes

OpenStudy (anonymous):

yes yes it was mistake by me i check it and i get 789 :P thanx very much for helping me do u need any thing in physics :P or math

OpenStudy (anonymous):

do u mean to say doubts?

OpenStudy (anonymous):

here you are.. Resistance = [resistivity] * [length] / [ cross sectional area] = [1.7*10^-8] * [ 0.2] / [ pi * (10^-3)^2 /4 ] = 4.33*10^-3 ohm Now, current = 3 / resistance = 3/ ( 4.33*10^-3) = 693 Amperes @Siddharth18 did I do sth wrong mate?

OpenStudy (anonymous):

@blacktigger how did you guys get 789 A?

OpenStudy (anonymous):

my resistance is coming out to be3.98*10^-3ohms

OpenStudy (anonymous):

when i checked my answer on the web site i found that 693 A is the right answer good job mate @bhaweshwebmaster

OpenStudy (anonymous):

@siddharth18 but u did agood joob also ty

OpenStudy (anonymous):

no big deal man,it was just the calculation. the approach is what matters more